Translate the following statement into an algebraic equation Let x represent the number.
6 times a number is 14 more than that number.

Answer by checkley71(8403)

6X+14=X
6X-X=-14
5X=-14
X=-14/5 ANSWER.
PROOF
6*-14/5+14=-14/5
-84/5+14=-14/5
(-84+70)/5=-14/5
-14/5=-14/5